# Attention-Guided Contrastive Role Representations for Multi-Agent Reinforcement Learning

### 08 Dec 2023 v1

@@ 25       Update the parameters of individual Q-network and the mixing network
In this paper, we use simple network structures for the trajectory encoder, the role encoder, and the attention mechanism. Specifically, the trajectory encoder contains a fully-connected multi-layer perceptron (MLP) and a GRU network with ReLU as the activation function, and encodes agent’s trajectory into a 128-dimensional embedding vector. The role encoder is a fully-connected MLP that transforms the 128-dimensional agent embedding into a 64-dimensional role representation. The setting of the mixing network is kept as the same as that of QMIX [14], where the architecture contains two 32-dimensional hidden layers with ReLU activation. Table 1 shows the details of network structures.
@@ Appendix C. Extended Experiments on Google Research Football
Fig. 7 presents the performance comparison between ACORM and baselines on three challenging GRF scenarios. It can be observed that QMIX obtains poor performance, since the tasks in GRF are more challenging than in the SMAC benchmark. In contrast, ACORM gains a significantly improved increase in the test win rate, especially in the first 1M training timesteps, which successfully demonstrates the effectiveness of our method evaluated on GRF benchmarks. Together with evaluations on SMAC domains, the same conclusion can still be drawn that ACORM outperforms baseline methods by a larger margin on harder tasks that demand a significantly higher degree of behavior diversity and coordination. In summary, experimental results on extended GRF environments are generally consistent with those on the SMAC benchmark.
@@ Appendix E. Hyperparameter Analysis on the Number of Clusters
Figure 11: Visualization with different numbers of clusters $K$ in negative pairs generation.

### 02 Mar 2024 v2

@@ 5 Conclusion and Discussion
Acknowledgements
The work was supported by the National Natural Science Foundation of China under Grant 62376122, Grant 62073160, Grant 62276126, and Grant 62176116.
@@ 25       Update the parameters of individual Q-network and the mixing network
In this paper, we use simple network structures for the trajectory encoder, the role encoder, and the attention mechanism. Specifically, the trajectory encoder contains a fully-connected multi-layer perceptron (MLP) and a GRU network with ReLU as the activation function, and encodes agent’s trajectory into a 128-dimensional embedding vector. The role encoder is a fully-connected MLP that transforms the 128-dimensional agent embedding into a 64-dimensional role representation. The setting of the mixing network is kept as the same as that of QMIX (Rashid et al., 2020), where the architecture contains two 32-dimensional hidden layers with ReLU activation. Table 1 shows the details of network structures.
@@ Appendix C. Extended Experiments on Google Research Football
Fig. 7 presents the performance comparison between ACORM and baselines on three challenging GRF scenarios. It can be observed that QMIX obtains poor performance, since the tasks in GRF are more challenging than in the SMAC benchmark. In contrast, ACORM gains a significantly improved increase in the test win rate, especially in the first 1M training timesteps, which successfully demonstrates the effectiveness of our method evaluated on GRF benchmarks. Together with evaluations on SMAC domains, the same conclusion can still be drawn that ACORM outperforms baseline methods by a larger margin on harder tasks that demand a significantly higher degree of behavior diversity and coordination. In summary, experimental results on extended GRF environments are generally consistent with those on the SMAC benchmark. .
Due to the very limited time for rebuttal revision, we only compare ACORM to QMIX and CDS, as other baselines (RODE, EOI, MACC, CIA) are not evaluated on GRF in their original papers. Also, we just show the performance within 2M training steps. We hope that the extended experimental evaluation could demonstrate adequate persuasiveness of our method. We are rushing to conduct the evaluation of baseline methods on GRF scenarios with more training timesteps, and trying to update them before the rebuttal deadline.
@@ Appendix E. Hyperparameter Analysis on the Number of Clusters
Figure 11: Visualization with different number of clusters $K$ in negative pairs generation.